OpenStudy (anonymous):

1) Change (8, 23°) into rectangular form. 2) Change (3, 2π) into rectangular form. 3) Change (5, 180°) into rectangular form.

hero (hero):

1. Easy 2. Easier 3. Easiest

OpenStudy (anonymous):

how about some answers then.

hero (hero):

Nah, I'm good. Giving answers don't help anyone learn anything.

OpenStudy (anonymous):

ok, then get off of my page and never comment again. @Hero

hero (hero):

Are you threatening me?

OpenStudy (anonymous):

good bye.

hero (hero):

Bye

jimthompson5910 (jim_thompson5910):

(8,23 degrees) = (r, theta) r = 8 theta = 23 degrees ------------------------------ x = r*cos(theta) x = 8*cos(23) x = ??? y = r*sin(theta) y = 8*sin(23) y = ??? The answer will be in the form (x,y)

OpenStudy (anonymous):

hmm..

jimthompson5910 (jim_thompson5910):

all you have to do is evaluate 8*cos(23) and 8*sin(23) using a calculator

OpenStudy (anonymous):

i go -0.5

OpenStudy (anonymous):

got*

jimthompson5910 (jim_thompson5910):

for which one, 8*cos(23) or 8*sin(23) ?

OpenStudy (anonymous):

DamnImBad

OpenStudy (anonymous):

do you know the answers?

OpenStudy (anonymous):

Nope, I don't remember how to do this

OpenStudy (anonymous):

@jim_thompson5910 gae u the fomula all you have to do is finish the problem

hero (hero):

@lalaloopsey, unblock me.

OpenStudy (anonymous):

answer my questions then..

OpenStudy (anonymous):

@Hero

hero (hero):

I'll see what I can do, but unblock me

hero (hero):

 r = 8